rootbeer003:

whats the inverse of y=9-x?

SkyVoltage43:

-x+y=9?

SkyVoltage43:

what lvl of math is this?

rootbeer003:

that's not a choice either y=9-x y=x-9 y=1/9-x y=1/x-9

rootbeer003:

do u know?

SkyVoltage43:

C

rootbeer003:

'c' Wanna tell me why?

SkyVoltage43:

nope

SkyVoltage43:

lol

rootbeer003:

Then get off my question

SkyVoltage43:

rude

rootbeer003:

So is your uncertain guess

SkyVoltage43:

im sure of my answer and you asked if i wanna tell you why its C

SkyVoltage43:

so i said no

SkyVoltage43:

i used a math website and it says its C so take the answer or leave cause no one else is gonna help

rootbeer003:

What website?

SkyVoltage43:

mathway

SkyVoltage43:

https://www.mathway.com/FiniteMath

rootbeer003:

ok thanks

SkyVoltage43:

welcome